OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y – 6 = −3x − 3 y – 6 = −3(x + 1) y = –3x + 3 3x + y = 3

hartnn (hartnn):

the point slope form : \(y-y_1=m(x-x_1)\)

hartnn (hartnn):

m= slope, x1,y1 is the point

OpenStudy (anonymous):

for point lope is given by y-y1=m(x-x1)

OpenStudy (anonymous):

here m=-3 and x1 =-1 and y=6

OpenStudy (anonymous):

thanks !!!